OpenStudy (anonymous):

simplify: 5x^3+10x^2+15x / 5x

OpenStudy (anonymous):

A. x2 + 10x2 + 15x B. 2x + 6 C. x2 + 2x + 3 D. x4 + 2x3 + 3x2

OpenStudy (anonymous):

\(\large \frac{5x^3+10x^2+15x}{5x}=\frac{5x^3}{5x}+\frac{10x^2}{5x}+\frac{15x}{5x} \) simplify that...., cancel common factors in each term...

OpenStudy (anonymous):

i don't know how to simplify fractions.

sam (.sam.):

Its basically just divide the top one with the bottom one

sam (.sam.):

From this,\[\frac{5x^3}{5x}\] 5 divided by 5 is? and x^3 divided by x is?

OpenStudy (anonymous):

5/5 = 1. I don't know how to do x^3/x